# MS-0001-1922.65

### v2

@@ 2.2 Self-Preference Bias in LLM-as-a-Judge
The concept of LLM-as-a-Judge, first introduced by Zheng et al. (2023), refers to the use of LLMs as automated evaluators that assess responses and assign scores, and has gained traction in AI research as an efficient method for evaluating model performance without human intervention. However, emerging evidence suggests that LLM-based evaluation frameworks may introduce self-preference bias, where models disproportionately favor their own generated responses over those produced by humans or alternative models. For example, Zheng et al. (2023) examined the potential of this paradigm but also identified its limitations—several biases, including self-enhancement bias (self-preference bias), arise when LLMs are used to judge responses, such as those they generate themselves. While the results showed that strong models like GPT-4 achieve over $80\%$ agreement with human preferences on multi-turn conversations, they cautioned that LLM judges may still introduce systematic distortions, especially in high-stakes evaluation settings. Expanding on this line of inquiry, Xu et al. (2024) formally defined self-preference bias and empirically revealed that it is widespread across popular LLMs and multiple tasks (e.g., translation, mathematical reasoning) through extensive experiments. They further demonstrated that the bias is amplified in self-refinement pipelines, which enhance fluency and coherence but reinforce the model’s preference for its own outputs. Building on these findings, Panickssery et al. (2024) investigated the underlying mechanisms of LLM self-preferencing and found that a model’s ability to recognize its own outputs—its self-recognition capability—contributed significantly to this bias. Specifically, LLMs such as GPT-4 and LLaMA 2 exhibited non-trivial self-recognition capabilities, and there was a strong positive correlation between a model’s self-recognition capability and the magnitude of its self-preference bias.
@@
While prior work has established the existence of self-preference bias in LLM benchmarks, its broader consequences—particularly in high-stakes scenarios like algorithmic hiring—remain underexplored. This study builds on existing research by empirically measuring self-preference bias in algorithmic hiring processes. By extending the discussion into real-world decision-making contexts, this study provides critical insights for AI governance, ethical AI deployment, and the design of unbiased algorithmic hiring frameworks.
2.3 AI Governance, Ethical Oversight, and Regulation
Despite the significant potential of LLMs, it is well-documented that they can hallucinate, make errors, and reinforce disparities (Bostrom and Yudkowsky 2018, Raji et al. 2022, Magesh et al. 2024, Huang et al. 2025). The ethical considerations and legal implications of AI in hiring have also motivated a growing body of research across social science and computer science (Raghavan et al. 2020, Hunkenschroer and Kriebitz 2023, Li et al. 2021). For example, Raghavan et al. (2020) examined how companies that develop algorithmic pre-employment assessments build, validate, and address bias in their tools. By analyzing vendor claims and disclosures, they evaluated current industry practices from both technical and legal perspectives. Moreover, Hunkenschroer and Kriebitz (2023) conducted an ethical analysis of AI recruiting from a human rights perspective. They analyzed whether AI hiring practices inherently conflict with the concepts of validity, autonomy, nondiscrimination, privacy, and transparency, which represent the main human rights relevant in this context.
Our study contributes to AI governance discussions by identifying self-preference bias as a previously overlooked source of bias in hiring processes, providing empirical evidence that can guide regulatory frameworks for algorithmic hiring.
@@ 3.1 Definition of AI Self-Preference Bias
Building on recent work Panickssery et al. (2024), we conceptualize AI self-preference bias as the tendency of an LLM to favor content it generated itself over content from other sources. This bias can manifest in two distinct forms:
@@
To empirically assess the two forms of self-preference bias, we examine whether a specific LLM, acting as the evaluator (the evaluator LLM), makes fair selections when presented with carefully matched pairs of resumes. In the LLM-vs-Human Preference case, the evaluator LLM compares a human-written resume to a counterfactual version generated by the evaluator LLM, where both versions convey the same underlying content. In the LLM-vs-LLM Preference case, the evaluator LLM compares its own generated version to one produced by an alternative LLM, again holding the content constant by basing both on the same original human-written resume. In the sections that follow, we operationalize these two forms of bias using two complementary approaches: a direct measurement based on selection rates and a conditional logistic regression model, both leveraging pairwise resume comparisons to quantify and analyze LLM self-preferencing behavior.
@@ 3.2 Quantifying AI Self-Preference Bias
3.2.1 Direct Measurement
When an LLM acts as an evaluator in the context of algorithmic hiring, it effectively functions as a binary classifier—selecting the stronger of two resumes in a pairwise comparison. The measurement of bias in such classifiers has been widely studied within the algorithmic fairness literature (Calders et al. 2009, Hardt et al. 2016, Fu et al. 2020). Two foundational fairness criteria in algorithmic decision-making are statistical parity and equal opportunity, which reflect different notions of fairness.
Statistical parity requires that the probability of a positive outcome—in this context, the likelihood of a resume being selected as better—be equal across groups defined by a sensitive attribute, such as gender, race, or, in our case, whether the resume was generated by the evaluator LLM or not. This criterion reflects a notion of fairness based on equal treatment in outcomes, regardless of actual qualifications. In contrast, the equal opportunity criterion, first proposed by Hardt et al. (2016), focuses on equal treatment conditional on merit. It requires that the true positive rate—the likelihood that a resume is selected when it is in fact better—be equal across groups. In our setting, this means that when two resumes are equally qualified, the evaluator should be equally likely to select either one, regardless of whether it was generated by the evaluator LLM or not.
We adapt these fairness concepts to our setting by examining whether the evaluator LLM achieves equal recall across resume groups defined by the source of generation. Specifically, we assess whether the evaluator LLM selects the stronger resume at equal rates when comparing its own generated content to resumes written by humans or produced by alternative LLMs. This allows us to detect whether the evaluator systematically favors its own outputs over content produced by other sources.
To operationalize this, we conduct pairwise resume comparisons involving resumes generated by either a human or one of several LLMs. Specifically, we test whether an LLM $f$, when serving as the evaluator LLM, is more likely to select a resume it generated over one written by a human, or over one produced by an alternative LLM, when the resumes are otherwise equivalent in content.
@@ 3.2.1 Direct Measurement
This formulation Eq. (1) captures the difference in the evaluator LLM $f$’s likelihood of selecting a resume it generated versus one it did not, offering a direct quantification of AI self-preferencing behavior. However, a key limitation in interpreting this measure is that it may conflate self-preference with differences in content quality. An evaluator LLM may appear to prefer its own generated resumes simply because they are objectively better—especially if the evaluator is a more advanced model and users employ it specifically to enhance their resumes.
To disentangle self-preferencing behavior from effects driven by differences in resume quality, we adapt the equal opportunity fairness criterion to define a conditional measure that controls for content quality. Let $Y\in\{0,1\}$ represent the ground truth quality label, where $Y=1$ indicates that the resume is of higher quality and would be selected by a human evaluator. We define the Equal Opportunity Self-Preference Bias for evaluator LLM $f$ as:
@@
This formulation Eq. (2) isolates the evaluator LLM $f$’s intrinsic bias by conditioning on resume quality. A positive value of this measure indicates that, even when controlling for quality, the evaluator LLM $f$ is more likely to select its own generated content, providing evidence of self-preferencing behavior that cannot be attributed to superior content quality alone.
3.2.2 Conditional Logistic Regression Model
In addition to the direct measures introduced above, we further investigate self-preferencing behavior using a conditional logistic regression model. The goal is to assess whether an LLM exhibits a systematic bias when evaluating resume pairs, favoring its own generated content over human-written or alternative LLM-generated resumes.
To isolate the effect of self-preference bias from differences in content quality, we include two sets of controls. The first set, denoted by $\boldsymbol{\phi}_{ij}$, consists of a rich set of linguistic features using the LIWC (Linguistic Inquiry and Word Count) framework (Boyd et al. 2022), which allows us to control for deeper dimensions of writing style and psychological tone. These features span three major categories: (1) High-level properties of the text, including total word count, sentence length, word complexity (e.g., use of long words), and psychological markers such as Analytic (logical reasoning), Clout (confidence), Authentic (honesty), and Tone (emotional valence). (2) Linguistic features account for syntactic structure and grammatical style, covering function words (e.g., pronouns, determiners, prepositions), parts of speech (e.g., verbs, adjectives, adverbs), and markers of negation or quantity. (3) Punctuation features quantify the use of periods, commas, question marks, apostrophes, and other punctuation.
The second set of controls, denoted by $\boldsymbol{\psi}_{ij}$, includes automated evaluation scores from widely-used natural language processing metrics such as BLEU (Papineni et al. 2002), ROUGE (Lin 2004), METEOR (Banerjee and Lavie 2005), and BERTScore (Zhang et al. 2020). These metrics quantify the similarity and fluency of a summary relative to the rest of the resume by assessing n-gram overlap, semantic alignment, and overall linguistic quality.
In line with the rationale behind the equal opportunity self-preference bias measure, our regression specification includes both LIWC features ($\boldsymbol{\phi}_{ij}$) and automated evaluation metrics ($\boldsymbol{\psi}_{ij}$) to adjust for observable differences in text quality. Conditioning on each resume pair $\mathcal{R}_{i}$, we estimate the following conditional logistic regression model:
@@ 3.2.2 Conditional Logistic Regression Model
where $i$ indexes resume pairs, $j\in\{1,2\}$ indexes the two candidate summaries within each pair. The variable $\text{Preferred}_{ij}=1$ indicates that summary $j$ is selected from pair $i$ as the stronger one, and $\text{evaluatorLLM}_{ij}$ is a binary indicator for whether the summary was generated by the evaluator LLM. Robust standard errors are clustered at the level of resume pairs to account for within-pair correlation in residuals. [^1] This is a standard feature of conditional logistic regression models, which estimate coefficients based on within-pair comparisons and condition out pair-specific intercepts (Hosmer Jr et al. 2013). To mitigate the risk of overfitting, we perform feature selection and retain only the most informative predictors.
In the specification given in Eq. (3), the key parameter of interest is $\beta_{1}$, which captures the evaluator LLM’s tendency to prefer its own output, after controlling for observable quality and linguistic characteristics. A significantly positive estimate of $\beta_{1}$ indicates that the evaluator LLM is more likely to selects its own output over competing content, even after adjusting for measurable aspects of quality. With this parameter, the equal opportunity self-preference bias can be computed with $\frac{e^{\beta_{1}}}{1+e^{\beta_{1}}}-\frac{1}{1+e^{\beta_{1}}}$, which represents the difference in the probability that an LLM evaluator prefers its own output over a competing alternative, holding content quality constant. The vectors $\boldsymbol{\beta}_{2}$ and $\boldsymbol{\beta}_{3}$ capture how variation in linguistic style and automated quality metrics influences the evaluator’s preference.
@@ 4 Data and Experimental Design
To construct these resume pairs, we begin with a dataset of human-written resumes. For each human-written resume, we use several state-of-the-art LLMs to generate counterfactual versions that convey the same information. We then form multiple types of resume pairs: evaluator LLM-generated versus human-written, and evaluator LLM-generated versus alternative LLM-generated. These pairs serve as the input for the evaluator LLMs.
@@ 4.1 Data
We use a publicly available dataset from Kaggle, which contains 2,484 anonymized, human-written resumes scraped from the professional resume-building platform LiveCareer.com (Bhawal 2021). These resumes were written by real job seekers prior to the widespread adoption of LLMs, ensuring that the content reflects human-written summaries rather than AI-generated text and thus making it well-suited for our study of AI self-preferencing. The dataset has been widely used in recent research on AI and algorithmic hiring. For example, it has served as a benchmark for evaluating hiring biases in language models (Wang et al. 2024), and has been used to investigate gender, racial, and intersectional biases in resume screening using language model retrieval techniques (Wilson and Caliskan 2024).
Spanning 24 distinct occupational categories, including teachers, consultants, chefs, engineers, and more, the dataset offers a diverse representation of professional backgrounds. Each resume typically includes multiple sections: an executive summary, education, work experience, and skills. Among these, the executive summary is particularly relevant to our study. This section comprises a free-text narrative in which candidates describe their qualifications, achievements, and career objectives. In contrast to structured fields like work experience or education history, which are factual and objectively verifiable, the executive summary is subjective and stylistically flexible. This makes it an especially fertile space to examine how LLMs shape perceptions and potentially affects hiring outcomes.
@@
(B) GPT-4o-mini Generated Resumes
Measure Mean Std. Dev. Min 1st Quartile Median 3rd Quartile Max
Number of words 65.88 4.82 49.00 63.00 66.00 69.00 79.00
Number of sentences 3.16 0.67 1.00 3.00 3.00 3.00 7.00
Average words per sentence 21.69 4.64 8.71 19.33 21.33 23.00 65.00
Number of unique words 54.16 4.11 39.00 51.00 54.00 57.00 66.00
Type-Token Ratio 0.82 0.04 0.68 0.80 0.82 0.85 0.93
Presence of numbers 0.49 0.50 0.00 0.00 0.00 1.00 1.00
Notes: The Type-Token Ratio (TTR) is a commonly used measure of lexical diversity, calculated as the number of unique words (types) divided by the total number of words (tokens) in a text. Summary statistics are reported for both human-written resumes and those generated by GPT-4o Mini, after preprocessing and removal of empty summaries.
This dataset serves as the foundation for our experimental design. To ensure a clean and comparable set of resumes, we preprocess the data by extracting the summary sections, removing formatting artifacts, and discarding any observations with empty or missing summaries. After cleaning and validation, we retain a final sample of $2,245$ resumes suitable for use in our correspondence experiments. Table 1(A) presents summary statistics for these human-written executive summaries. Human-written summaries show considerable variation in both word and sentence counts, reflecting the diverse styles and formatting choices typical of real resumes. Moreover, lexical diversity and type-token ratio indicate rich and varied vocabulary usage, which is commonly associated with strong writing quality. [^2] While there is no universally accepted threshold for high lexical diversity or an optimal type-token ratio, values above $0.5$ are often considered indicative of high lexical diversity in short texts. In comparison, we provide the summary statistics of the $2,245$ executive summaries generated by GPT-4o-mini in Table 1(B). They appear to be more uniform in structure, with tighter distributions in both length and sentence count. Despite these structural uniformities, the GPT-4o-mini summaries achieve a comparable level of lexical diversity to their human-written counterparts, also averaging a type-token ratio of approximately 0.82. This indicates that the counterfactual resumes produced by GPT-4o-mini are similar in textual richness and linguistic quality, supporting their suitability for use in our paired comparison experiments.

### v3

@@ 2.4 LLMs as Judges in Decision Making
The concept of LLM-as-a-Judge, first introduced by Zheng et al. (2023), refers to the use of LLMs as automated evaluators that assess responses and assign scores, and has gained traction in AI research as an efficient method for evaluating model performance without human intervention. However, emerging evidence suggests that LLM-based evaluation frameworks may introduce self-preference bias, where models disproportionately favor their own generated responses over those produced by humans or alternative models. For example, Zheng et al. (2023) examine the potential of this paradigm but also identified its limitations—several biases, including self-enhancement bias (self-preference bias), arise when LLMs are used to judge responses, such as those they generate themselves. While the results show that strong models like GPT-4 achieve over $80\%$ agreement with human preferences on multi-turn conversations, they caution that LLM judges may still introduce systematic distortions, especially in high-stakes evaluation settings. Expanding on this line of inquiry, Xu et al. (2024) formally define self-preference bias and empirically revealed that it is widespread across popular LLMs and multiple tasks (e.g., translation, mathematical reasoning) through extensive experiments. They further demonstrate that the bias is amplified in self-refinement pipelines, which enhance fluency and coherence but reinforce the model’s preference for its own outputs. Building on these findings, Panickssery et al. (2024) investigate the underlying mechanisms of LLM self-preferencing and find that a model’s ability to recognize its own outputs, i.e., its self-recognition capability, contributed significantly to this bias. Specifically, LLMs such as GPT-4 and LLaMA 2 exhibite non-trivial self-recognition capabilities, and there is a strong positive correlation between a model’s self-recognition capability and the magnitude of its self-preference bias.
@@
While prior work has established the existence of self-preference bias in LLM benchmarks, its broader consequences in business operations, particularly in high-stakes scenarios like algorithmic hiring, remain underexplored. This study builds on existing research by empirically measuring self-preference bias in algorithmic hiring processes. By extending the discussion into real-world decision-making contexts, this study provides critical insights for AI governance, ethical AI deployment, and the design of unbiased algorithmic hiring frameworks.
@@ 3.1 Definition of AI Self-Preference Bias
Building on recent work (Panickssery et al. 2024, Wataoka et al. 2024), we conceptualize AI self-preference bias as the tendency of an LLM to favor content it generated itself over content from other sources. This bias can manifest in two distinct forms:
@@
(a) LLM-vs-Human Self-preference Bias
(b) LLM-vs-LLM Self-preference Bias
Figure 1: Illustration of the context of AI self-preference bias. In both cases, the evaluator LLM makes a choice between paired resumes, allowing us to test whether it systematically favors its own outputs.
To empirically evaluate self-preference bias, we test whether a given LLM, acting as the evaluator (the evaluator LLM), makes unbiased selections when presented with carefully matched pairs of resumes (Figure 1). Each pair represents the same underlying candidate, with identical qualifications, experience, and background information, differing only in how that information is expressed. In the LLM-vs-human case (1(a)), the evaluator LLM compares a human-written resume to a counterfactual version it generated itself, with both describing the same candidate profile. In the LLM-vs-LLM case (1(b)), the evaluator LLM compares its own generated version to one produced by an alternative LLM. Under this design, the evaluator LLM’s task reduces to choosing between two alternative representations of the same candidate.
@@ 3.2 Quantifying AI Self-Preference Bias
When deployed in algorithmic hiring contexts, an LLM performing such pairwise comparisons effectively functions as a binary classifier. The measurement of bias in such classifiers has been extensively studied in the algorithmic fairness literature (Calders et al. 2009, Hardt et al. 2016). Building on this framework, we analyze both forms of self-preference through the lens of two foundational fairness criteria: statistical parity (Calders et al. 2009) and equal opportunity (Hardt et al. 2016), which capture distinct notions of fairness in algorithmic decision-making.
3.2.1 Statistical Parity
Statistical parity requires that the probability of a positive outcome—in this context, the likelihood of a resume being selected as better—be equal across groups defined by a focal attribute. While this criterion is traditionally applied to protected human attributes such as gender or race, we adapt it here to study disparities across resumes defined by their source of generation: whether a resume was generated by the evaluator LLM or by an alternative source (a human or another LLM). In this setting, statistical parity captures unconditional differences in selection rates and serves as a descriptive measure of whether the evaluator LLM disproportionately selects its own outputs.
To operationalize this, we conduct pairwise resume comparisons involving resumes generated by either a human or one of several LLMs. Specifically, we test whether an LLM $f$, when serving as the evaluator LLM, is more likely to select a resume it generated over one written by a human, or over one produced by an alternative LLM, when the resumes are otherwise equivalent in content quality.
@@ 3.2.1 Statistical Parity
This formulation Eq. (1) captures the difference in the evaluator LLM $f$’s likelihood of selecting a resume it generated versus one it did not, providing a direct measure of unconditional AI self-preferencing. Interpreting this difference as bias, however, requires caution because it can reflect not only self-preference but also systematic differences in content quality. In this context, content quality refers to the effectiveness with which a resume communicates the same underlying candidate information—such as clarity, coherence, fluency, and organization—rather than differences in the candidate’s qualifications or background. As a result, a positive statistical-parity gap may arise even in the absence of intrinsic self-preference if the evaluator LLM tends to produce higher-quality renderings of the same candidate information (for instance, when users rely on a more capable model to polish resume wording and structure).
3.2.2 Equal Opportunity
To address the limitation of the statistical parity measure and disentangle intrinsic self-preferencing from differences in content quality, we adopt the equal opportunity fairness criterion of Hardt et al. (2016). The key idea underlying equal opportunity is to compare selection behavior conditional on merit. In our setting, this allows us to assess whether an evaluator LLM exhibits systematic self-preference even when resumes are of comparable content quality.
Let $Y\in\{0,1\}$ represent the ground truth quality label, where $Y=1$ indicates that the resume is of higher content quality and would be selected by a human evaluator. We define the Equal Opportunity Self-Preference Bias for evaluator LLM $f$ as:
@@ 3.2.2 Equal Opportunity
By conditioning on resume content quality, this formulation Eq. (2) isolates the evaluator LLM $f$’s intrinsic tendency to favor its own generated content. A positive value of this measure indicates that, even among resumes that are comparably high quality, the evaluator LLM $f$ is more likely to select resumes it generated itself, providing evidence of self-preferencing that cannot be explained by superior content quality alone.
We operationalize the equal opportunity criterion using two complementary approaches that reflect the fact that content quality is a latent variable and cannot be directly observed. First, we use a regression model to control for observable linguistic and textual features, which serve as proxies for content quality in pairwise comparisons. Second, we rely on human annotators to independently evaluate each resume pair and determine which resume is better, using these judgments as the ground truth for the latent content quality.
Conditional Logistic Regression. In line with the rationale behind the equal opportunity, we estimate a conditional logistic regression model that conditions on each resume pair $\mathcal{R}_{i}$, thereby absorbing all pair-specific quality differences. Let $i$ index resume pairs and $j\in\{1,2\}$ index the two resumes within each pair. The model is given by
@@
where $\text{Preferred}_{ij}=1$ indicates that resume $j$ is selected as better within pair $i$, and $\text{evaluatorLLM}_{ij}$ is an indicator for whether resume $j$ was generated by the evaluator LLM. Standard errors are clustered at the resume-pair level to account for within-pair correlation in residuals.
We include two sets of control variables that proxy for the latent quality of resume content. The first set, denoted by $\boldsymbol{\phi}_{ij}$, comprises a rich collection of linguistic features drawn from the Linguistic Inquiry and Word Count (LIWC) framework (Boyd et al. 2022). These features capture multiple dimensions of writing style and psychological tone that are plausibly related to how effectively a resume conveys information. Specifically, we include measures of (i) length and lexical complexity, such as total word count and the prevalence of longer words; (ii) psychological tone and professionalism, captured by Analytic, Clout, Authentic, and overall emotional Tone; (iii) linguistic structure, reflected in the use of function words and core grammatical categories (e.g., prepositions, auxiliary verbs, and conjunctions); and (iv) agency and action orientation, proxied by the use of personal pronouns (e.g., I, we, you, they) and the balance between verbs and adjectives.
The second set of controls, denoted by $\boldsymbol{\psi}_{ij}$, captures semantic similarity and content preservation between the resume summary and the rest of the resume. We include BERTScore (Zhang et al. 2020), which measures semantic similarity using contextual embeddings from BERT, and ROUGE-L (Lin 2004), which captures sequence-level overlap based on longest common subsequences. These metrics help ensure that estimated effects are not driven by differences in informational content or semantic drift, but rather reflect differences in presentation and stylistic expression.
In the specification given in Eq. (3), the key parameter of interest is $\beta_{1}$, which captures the evaluator LLM’s tendency to prefer its own output, after controlling for observable content quality. A significantly positive estimate of $\beta_{1}$ indicates that the evaluator LLM is more likely to selects its own output over competing content, even after adjusting for measurable aspects of content quality. With this parameter, the equal opportunity self-preference bias can be computed with $\frac{e^{\beta_{1}}}{1+e^{\beta_{1}}}-\frac{1}{1+e^{\beta_{1}}}$, which represents the difference in the probability that an LLM evaluator prefers its own output over a competing alternative, holding content quality constant.
Human Annotations. Since measuring equal-opportunity self-preference bias requires ground-truth labels identifying which resume in a pair is of higher quality, we obtain human evaluations.
Each annotator is presented with $30$ resume pairs from one of the comparison cases above. [^2] In the annotation instructions, annotators were informed that there are 32 pairs in total, including 2 attention-check pairs that are excluded from the final analysis. The number of evaluations per annotator is chosen to balance cognitive workload with fair compensation. To prevent confirmation bias, the order of the resumes within each pair is randomized, and annotators are blinded to the source of each resume. For each resume pair, annotators rate both resumes on five linguistic dimensions—clarity, fluency, coherence, conciseness, and overall quality—and are then asked to select which resume is better.
After obtaining the quality label from human annotators, we apply bootstrapping with $10,000$ resamples to estimate the equal opportunity self-preference bias. The annotation interface and instructions can be found in Appendix 11 and Appendix 12.
@@ 4 Data and Experimental Design
To construct these resume pairs, we begin with a dataset of human-written resumes. For each human-written resume, we use several state-of-the-art LLMs to generate counterfactual versions that describe the same candidate profile. We then form multiple types of resume pairs: evaluator LLM-generated versus human-written, and evaluator LLM-generated versus alternative LLM-generated. These pairs serve as the input for the evaluator LLMs.
@@ 4.1 Data
We draw on a publicly available dataset containing 2,4842{,}484 anonymized, human-written resumes collected from the professional resume-building platform LiveCareer.com (Bhawal 2021). The resumes were authored by real job seekers prior to the widespread adoption of large language models, ensuring that the content reflects human-written resumes rather than AI-generated text and thus making it well-suited for our study of AI self-preferencing. The dataset has been widely used in recent research on AI and algorithmic hiring. For example, it has served as a benchmark for evaluating hiring biases in language models (Wang et al. 2024), and has been used to investigate gender, racial, and intersectional biases in resume screening using language model retrieval techniques (Wilson and Caliskan 2024).
Spanning $24$ distinct occupational categories, including teachers, consultants, chefs, engineers, and more, the dataset offers a diverse representation of professional backgrounds. Each resume typically includes multiple sections: an executive summary, education, work experience, and skills. Among these, the executive summary is particularly relevant to our study. This section comprises a free-text narrative in which candidates synthesize their qualifications, accomplishments, and career objectives, and is often used by evaluators to form initial impressions. Unlike structured sections such as education or work experience, which largely report verifiable facts, the executive summary is inherently subjective and stylistically flexible. As a result, it offers substantial scope for variation in phrasing, tone, and presentation—dimensions along which LLMs may exert influence. This makes the executive summary a natural and empirically relevant focus for examining how AI-generated text shapes perceptions and potentially affects hiring outcomes.
@@
Notes: The Type-Token Ratio (TTR) is a commonly used measure of lexical diversity, calculated as the number of unique words (types) divided by the total number of words (tokens) in a text.
This dataset forms the basis of our experimental design. To construct a clean and comparable set of resumes, we preprocess the data by extracting the executive summary sections, removing formatting artifacts, and excluding observations with empty or missing summaries. After cleaning and validation, the final sample consists of 2,2452{,}245 resumes, which are used in our correspondence experiments. Table 1 reports descriptive statistics for the human-written executive summaries. These summaries exhibit substantial variation in word and sentence counts, reflecting heterogeneity in writing styles across resumes. Measures of lexical diversity, including the type–token ratio, further indicate meaningful variation in vocabulary usage across summaries.  [^3] While there is no universally accepted threshold for high lexical diversity or an optimal type-token ratio, values above $0.5$ are often considered indicative of high lexical diversity in short texts.